OpenStudy (anonymous):

If a projectile is fired from the origin (on earth), with an initial velocity v0, and at an angle theta above the horizontal, how much time does it take to cross a line starting at the origin at an angle alpha < theta above the horizontal

OpenStudy (anonymous):

|consider the line as an inclined plane with a slope alpha.from the basement a body is thrown in an angle (theta-alpha).Now the gravitational acceleration acting on the body is (taking the slope as the ground) gCosalpha in the vertical direction and gSinalpha in the horigental direction. time to reach highest point will be \[V \cos( \theta -\alpha)divg time \cos \alpha =t\] so 2t is the time reqired dw:1413139397971:dw|

OpenStudy (anonymous):

Figured out that question. Working on figuring out your formula. Is that supposed to be \[v \cos(\theta-\alpha)\div(g \cos(\alpha)) = t\]

OpenStudy (anonymous):

And thank you

OpenStudy (anonymous):

I see where that formula comes from, so that should be it.